can you please help me answer this equation.
3/x+1 + 2/x-1=

Answer by RAY100(1637)

3/(x+1) +2/(x-1)
.
LCD = (x+1)(x-1) = x^2 -1
.
{ 3(x-1)+2(x+1) }/ { (x+1)(x-1)}
.
{ 3x-3+2x+2 } / {x^2 -1}
.
{5x -1} / {x^2 -1}
.
check, Let x=2,original and answer = 3
.